The ratio of difference int he wavelengths of 1st and 2nd lines of lyman seres in H-like atim of difference in wavelength for 2nd and 3rd lines of same series is -
  • 2.5:1
  • 3.5:1
  • 4,5:1
  • 5.5:1

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
1/λ = 1.097 x 10^7 ( 1 - 1/n^2)

n=2 =>
1/λ = 1.097 x 10^7(1 - 1/4) = 0.82275 x 10^7 per m
=> λ = 1.215 x 10^(-7) m ------------eq(1)

n=3 =>
1/λ = 1.097 x 10^7(1 - 1/9) = 0.9751 x 10^7 per m
=> λ = 1.026 x 10^(-7) m …......................eq(2)

n=4 =>
1/λ = 1.097 x 10^7(1 - 1/16) = 1.0284 x 10^7 per m
=> λ = 0.972 x 10^(-7) m...............eq(3)

ratio is ew 1-2/2-3
=3.5:1